OpenStudy (anonymous):

Between 7:00 AM and 8:00 AM, trains arrive at a subway station at a rate of 15 trains per hour per minute). The following formula from statistics can be used to determine the probability that a train will arrive within t minutes of 7:00 AM. F(t) = 1 - e-0.25t Determine how many minutes are needed for the probability to reach 20%.

OpenStudy (anonymous):

F(t) = 1 - e - 0.25t Let t = 20% = 0.2 F(0.2) = 1 - e - 0.25(0.2) = 1 - 0.05 - e = 0.95 - e